OpenStudy (samsan9):

x+3 is a factor of P(x)=x^3-7x^2+15x-9. true or false

jimthompson5910 (jim_thompson5910):

if x-k is a factor of p(x), then p(k) = 0

jimthompson5910 (jim_thompson5910):

in this case, k = -3 since x+3 = x-(-3) = x-k

jimthompson5910 (jim_thompson5910):

P(-3) = ???

OpenStudy (samsan9):

so you just have to plug in -3 and if it doesn'tequal zero then it isn't a factor?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

http://www.purplemath.com/modules/factrthm.htm